the trip is 80 miles long, 40 miles out and 40 miles back. If the bus travels at a speed of 50 mph on the way out and 30 mph on the wat back. How many minutes will the class be on the bus?
vonda.walker@verizon.net

Answer by nerdybill(7384)

the trip is 80 miles long, 40 miles out and 40 miles back. If the bus travels at a speed of 50 mph on the way out and 30 mph on the wat back. How many minutes will the class be on the bus?
.
Apply the distance formula:
d = rt
but in your case you're interested in time so solving for t:
t = d/r
.
Time going out:
t = d/r
t = 40/50 = 4/5 hour
Time coming back:
t = d/r
t = 40/30 = 4/3 hour
.
Total = 4/5 + 4/3
Total = 12/15 + 20/15
Total = 22/15 hour
.
Convert "total time" to minutes:
22/15 * 60 = 88 minutes